# The Universal Weight Subspace Hypothesis

### v1

@@ 1 Introduction
Several lines of prior research have hinted at phenomena consistent with our joint (universal) subspace hypothesis. For example, Neural Tangent Kernel (NTK) theory demonstrates that, in the infinite-width limit, the training dynamics of deep networks are governed by a kernel largely invariant to task specifics (jacot2018ntk). Similarly, research into mechanistic interpretability—specifically its own universality hypothesis (olah2020zoom; chughtai2023toymodeluniversalityreverse)—has uncovered recurring circuits and patterns within specific layers of toy or vision networks, lending indirect support to the concept of universality. Other phenomena, such as the lottery ticket hypothesis (frankle2019lottery) and mode connectivity (garipov2018mode), provide further evidence for the existence of reusable, low-dimensional structures in neural networks. notably, Krizhevsky2012ImageNet observed that the first layer of convolutional networks consistently learns Gabor-like filters across diverse vision tasks. More recently, studies by guth2024universalityneuralencodingscnns; guth24; kaushik2025eigenloraxrecyclingadaptersprincipal have shown initial evidence of recurring eigenvectors in certain layers of convolutional neural networks trained on natural images. While works such as martin2025setolsemiempiricaltheorydeep; Mao_2024 have explored neural properties to explain generalization or loss landscape convergence, they do not address the convergence of parametric properties across distinct models trained on disjoint data. Importantly, in contrast to the abstract, weak, or speculative notions of universality presented in earlier works like chughtai2023toymodeluniversalityreverse; olah2020zoom - which focus primarily on representations and are arguably easier to demonstrate due to direct data dependency - ours is the first work to provide concrete evidence and a clear universal hypothesis at the neural parameter or weights level.
In our analysis, we present compelling empirical evidence for the existence of universal subspaces within LoRA adapters across different modalities and tasks. We initially focus on LoRA adapters due to their ease of training and the ability to collect a large number of adapters for diverse tasks, models, and datasets, which enables robust evaluation of our hypothesis. E.g., we demonstrate the emergence of a universal subspace across approximately 500 LoRA adapters for the Mistral-7B (jiang2023mistral7b) model. We further extend our investigation to the full weight space, where we observe similar universality, extracting sparse, low-rank universal subspaces from about 500 Vision Transformer models and 50 LLaMA3-8B models, each trained on different datasets and initializations.
@@
To date, our work presents the most rigorous empirical evidence for the existence of “Universality” within the parameter space of deep neural networks. This geometric universality offers a novel vantage point for investigating fundamental neural properties, including generalization, grokking, catastrophic forgetting, and dataset efficiency. Crucially, the pervasive nature of this universal property underscores the relative primacy of model architecture over other factors in shaping the learned parameter space (.) However, our findings also delineate clear frontiers for future inquiry. We leave open the question of cross-architectural comparison: how do the universal subspaces of distinct architectures differ, and can we explicitly design architectures to optimize the geometry of this subspace? Furthermore, a fundamental question emerges from our results regarding the implications of convergence: if neural networks systematically collapse into the same subspace - thereby inheriting shared biases, capabilities, and failure modes - is this lack of diversity a fundamental bottleneck, and should we develop methods specifically designed to break this convergence?
@@ 2 Notations, Definitions and Theoretical Analysis
Our theoretical analysis models predictors as elements of a Hilbert space, for example a reproducing kernel Hilbert space (RKHS), while our experiments are conducted with practical large-scale models such as transformers and LoRA-based variants. Modeling predictors in a Hilbert space (kernel) framework is standard when analyzing aspects such as generalization and inductive bias of modern deep architectures, and has been widely used to approximate or interpret the behavior of large neural networks in practice (ortiz-jimenez2023taskarithmetic; wei2019regularization; chen2021deep; belfer2024spectral; bietti2019kernelperspectiveregularizingdeep). We aim to understand whether the shared structure across tasks can be consistently recovered from data as number of tasks increase. Specifically, each task has an associated ground-truth predictor $f_{t}^{\star}$, and we are interested in the covariance (second-moment) operator $\mathcal{S}$ that captures the common subspace spanned by these predictors. Since in practice we only observe finite samples per task and learn approximate predictors $\hat{f}_{t}$, two sources of error arise: (i) variability due to having finitely many tasks, and (ii) estimation noise within each task. Our goal is to establish conditions under which the empirical operators built from $\hat{f}_{t}$ concentrate around $\mathcal{S}$, and to show that the learned top-$k$ subspace converges to the true one, with convergence rates that separately reflect the number of tasks and the accuracy of per-task learning.
@@
Here $\mathcal{R}_{n_{t},D_{t}}(\mathcal{H})$ represents Rademacher complexity of the solutions within Hilbert space $\mathcal{H}$ over $n_{t}$ samples drawn i.i.d. from $D_{t}$ This form is satisfied, for example, by strongly convex regularized ERM in an RKHS (e.g., kernel ridge regression or NTK ridge), under bounded kernel norm and sub-Gaussian response noise (barlette2003rademacher).
@@ 3.1 Analysis methodology
Since there is no current method that enables us to compare subspaces of models with different architectures, we focus on large number of models trained on the same architecture. To this end, we perform analysis using Low rank adapters (hu2021lora) (LoRA) as well as classical weights of transformer and CNN (Convolutional Neural Network) architectures. For all our experiments, unless stated otherwise, we perform Order 1-2 HOSVD only, to ensure that our methodology works even in the simplest case. Algorithm˜1 provides the algorithm we implement. Refer to Appendix˜B for discussion regarding secondary subspace and how to choose the number of top components.
@@ 3.2.1 Lower-rank joint subspaces in CNNs, LoRA and Finetuned models
In smaller and conventional architectures such as CNNs, evidence for universal structure has been more limited but suggestive. Early work observed that the first convolutional layer often learns Gabor-like filters across diverse vision tasks (Krizhevsky2012ImageNet). More recently, works report recurring eigenvectors in certain CNN layers trained on natural images (guth24; guth2024universalityneuralencodingscnns).
@@
In order to conduct a more real-world experiment, we choose to run the subspace analysis for LoRA hu2021lora models simply because they are available in abundance in public domain. Given LoRA models distinctly capture task specific directions as they show weak alignment with the original weights hu2021lora, they form a good main model parameter alternative to run our subspace analysis and verify whether this holds true. We spectrally decompose (Section 3.1) LoRA’s submatrices individually, each concatenated across all the available finetuned LoRAs and choose top $k$ spectral basis. This setup allows us to truly stress test the Universal Subspace.
@@
In our first experiment, we use LoRA models trained on 500 natural instruction tasks (wang-etal-2022-super) using Mistral-7B-Instruct-v0.2 (jiang2023mistral7b) as the base (brüelgabrielsson2024compressserveservingthousands). Each LoRA, individually, is at least of rank 16. Figure 3 shows the results of our analysis. 3(a) shows the subspace analysis of individual layers of the Mistral model. Each bar corresponds to the Eigenvalue or explained variance of every unique component (unique for all models, trained in disjoint/different datasets), showing that the parameters of all 500 models can be well approximated by a finite low-rank subspace for all layers. Note that for visualization purposes, the graphs only show the top 100 components, but there exists a significantly larger number of basis. 3(b) provides a summary of this subspace analysis, (summarized) for all of the layers of the 500 Mistral models, further validating the idea of a layer-wise universal subspace. Please refer to the appendix for all other and larger versions of the given spectral plots.
@@
We extend our analysis to text-to-image generation using Stable Diffusion-XL (sdxl). A universal subspace is extracted from publicly available LoRAs on HuggingFace (von-platen-etal-2022-diffusers). When projecting individual LoRAs into this subspace, the resulting generations preserve visual quality and style (Figure 5). CLIP-based evaluations (Table 1) show that the universal subspace even outperforms individual LoRAs in some cases, possibly due to denoising effects previously observed in (sharma_laser_2023).
@@
In order to test the ability of condensing many models into a single universal subspace, we compare our method with SOTA model merging/combination methods in Table 2. We compare our universal subspace inspired combination approach against six state-of-the-art, gradient-free baselines: RegMean (jin2023dataless), Task Arithmetic (TA) (ilharco2023editing), TIES (yadav2023tiesmerging), DARE-TIES (YuDare), KnOTS-TIES, and KnOTS-DARE-TIES (knots). RegMean aligns task-specific updates by solving a layer-wise linear regression problem, requiring transformation matrices for each model. TA merges models by linearly combining parameters, but relies on tuning scaling coefficients on a validation set for optimal performance. TIES extends TA with magnitude-based pruning and sign conflict resolution, introducing additional hyperparameters such as pruning thresholds, while DARE-TIES combines random Bernoulli pruning with TIES’ sign resolution, also requiring tuning of pruning probability. KnOTS-TIES and KnOTS-DARE-TIES further apply SVD-based subspace alignment before merging, but still inherit the need for coefficient or pruning hyperparameter selection. In contrast, our universal subspace method, analytically computes the merging coefficients based solely on the geometry of a shared, low-rank universal subspace identified across models, requiring no iterative tuning or validation data-although optional finetuning is possible if data is available. Furthermore, because our subspace is intrinsically low-rank, the merged model contains significantly fewer parameters than any individual models, offering both computational efficiency and theoretical alignment guarantees not present in the baselines. Empirically, our approach achieves higher average accuracy (see Table 2), while reducing parameter count, thus enabling scalable and robust model merging without heuristic pruning or validation overhead. We note that we did not optimize our merging process and better results nearing finetuned performance may be achieved.
@@ 3.3 Low rank shared universal subspaces in classical weights
We further extend this analysis to 50 finetuned LLaMA3-8B models, 177 GPT-2 models, and Flan-T5 models (trained on GLUE glue datasets) again sourced from HuggingFace without filtering. As shown in Figure 6, a small number of directions capture dominant structure across models spanning diverse and distinct datasets and tasks. More details are provided in the Section˜B.2. This is, to our knowledge, the first instance of compressing such a large and diverse collection of foundation models into a unified subspace, highlighting its potential for large-scale model reuse and environmental efficiency.
@@ 3.3.1 Finding universal subspaces and applying them to future tasks
We present two experiments - Image Classification using ViT-base and Natural Language Understanding using GLUE benchmark glue with RoBERTa-base model. Both involve creating a universal subspace using publicly available LoRA adapters. Details are provided in the Appendix˜C. For the GLUE benchmark, we follow the same setup as  (kopiczko_vera_2023) considering the 6 tasks - CoLA, MRPC, SST-2, QNLI, RTE and STS-B while omitting the time-intensive MNLI and QQP tasks. We initialize our universal subspace using a leave-one-out-setup, where the subspace is calculated using components of all but one LoRA adapter for which the coefficients are learned. For image classification, we utilize publicly available ViT LoRAs to extract our universal subspaces taking care that the data any of these pretrained LoRAs have not seen the data we will be training our coefficients on.
@@ 4 Discussion
Why do these universal subspaces emerge? While the precise mechanisms driving this phenomenon remain an open area of investigation, several theoretical factors likely contribute to the emergence of these shared structures. First, neural networks are known to exhibit a spectral bias toward low-frequency functions, creating a polynomial decay in eigenvalues that concentrates learning dynamics into a small number of dominant directions (belfer2024spectral; bietti2019kernelperspectiveregularizingdeep). Second, modern architectures impose strong inductive biases that constrain the solution space: convolutional structures inherently favor local, Gabor-like patterns (Krizhevsky2012ImageNet; guth24), while attention mechanisms prioritize recurring relational circuits (olah2020zoom; chughtai2023toymodeluniversalityreverse). Third, the ubiquity of gradient-based optimization – governed by kernels that are largely invariant to task specifics in the infinite-width limit (jacot2018ntk) – inherently prefers smooth solutions, channeling diverse learning trajectories toward shared geometric manifolds (garipov2018mode). If these hypotheses hold, the universal subspace likely captures fundamental computational patterns that transcend specific tasks, potentially explaining the efficacy of transfer learning and why diverse problems often benefit from similar architectural modifications.
@@ 5 Limitations and Future Work
Although we provide conclusive results towards the existence and utility of universal shared subspaces, the current analysis has scope for future research, such as limited interpretability of the shared subspace and the corresponding directions. While it is a critical area of research, it is extremely cumbersome to demonstrate interpretability of the principal directions for each layer of the network. To the best of our knowledge we are not aware of any other literature that performs such an in-depth analysis of the weight space of large models across diverse tasks, data modalities and model architectures. The current approach to approximating a universal subspace relies on pretrained task-specific models (predictors) for tasks, which may not be readily available for new tasks. An interesting direction for future research would be to explore model independent methods for learning a universal shared subspace, potentially derived directly from data. Furthermore, the conditions proposed in ortiz-jimenez2023taskarithmetic for enabling task arithmetic rely on localized eigenfunctions which are not conducive to learning a shared universal subspace. As a result, performing task arithmetic within the current framework of a shared universal subspace is non-trivial and warrants further investigation.

### v2

@@ 1 Introduction
Several lines of prior research have hinted at phenomena consistent with our joint (universal) subspace hypothesis. For example, Neural Tangent Kernel (NTK) theory demonstrates that, in the infinite-width limit, the training dynamics of deep networks are governed by a kernel largely invariant to task specifics (Jacot et al., 2018). Similarly, research into mechanistic interpretability—specifically its own universality hypothesis (Olah et al., 2020; Chughtai et al., 2023)—has uncovered recurring circuits and patterns within specific layers of toy or vision networks, lending indirect support to the concept of universality. Other phenomena, such as the lottery ticket hypothesis (Frankle & Carbin, 2019) and mode connectivity (Garipov et al., 2018), provide further evidence for the existence of reusable, low-dimensional structures in neural networks. notably, Krizhevsky et al. (2012) observed that the first layer of convolutional networks consistently learns Gabor-like filters across diverse vision tasks. More recently, studies by Guth & Ménard (2024); Guth et al. (2024); Kaushik et al. (2025) have shown initial evidence of recurring eigenvectors in certain layers of convolutional neural networks trained on natural images. While works such as Martin & Hinrichs (2025); Mao et al. (2024) have explored neural properties to explain generalization or loss landscape convergence, they do not address the convergence of parametric properties across distinct models trained on disjoint data. Importantly, in contrast to the abstract, weak, or speculative notions of universality presented in earlier works like Chughtai et al. (2023); Olah et al. (2020) - which focus primarily on representations and are arguably easier to demonstrate due to direct data dependency - ours is the first work to provide concrete evidence and a clear universal hypothesis at the neural parameter or weights level.
In our analysis, we present compelling empirical evidence for the existence of universal subspaces within LoRA adapters across different modalities and tasks. We initially focus on LoRA adapters due to their ease of training and the ability to collect a large number of adapters for diverse tasks, models, and datasets, which enables robust evaluation of our hypothesis. E.g., we demonstrate the emergence of a universal subspace across approximately 500 LoRA adapters for the Mistral-7B (Jiang et al., 2023) model. We further extend our investigation to the full weight space, where we observe similar universality, extracting sparse, low-rank universal subspaces from about 500 Vision Transformer models and 50 LLaMA3-8B models, each trained on different datasets and initializations.
@@
To date, our work presents the most rigorous empirical evidence for the existence of “Universality” within the parameter space of deep neural networks. This geometric universality offers a novel vantage point for investigating fundamental neural properties, including generalization, grokking, catastrophic forgetting, and dataset efficiency. Crucially, the pervasive nature of this universal property underscores the relative primacy of model architecture over other factors in shaping the learned parameter space. However, our findings also delineate clear frontiers for future inquiry. We leave open the question of cross-architectural comparison: how do the universal subspaces of distinct architectures differ, and can we explicitly design architectures to optimize the geometry of this subspace? Furthermore, a fundamental question emerges from our results regarding the implications of convergence: if neural networks systematically collapse into the same subspace - thereby inheriting shared biases, capabilities, and failure modes - is this lack of diversity a fundamental bottleneck, and should we develop methods specifically designed to break this convergence?
@@ 2 Notations, Definitions and Theoretical Analysis
Our theoretical analysis models predictors as elements of a Hilbert space, for example a reproducing kernel Hilbert space (RKHS), while our experiments are conducted with practical large-scale models such as transformers and LoRA-based variants. Modeling predictors in a Hilbert space (kernel) framework is standard when analyzing aspects such as generalization and inductive bias of modern deep architectures, and has been widely used to approximate or interpret the behavior of large neural networks in practice (Ortiz-Jimenez et al., 2023; Wei et al., 2019; Chen & Xu, 2021; Belfer et al., 2024; Bietti et al., 2019). We aim to understand whether the shared structure across tasks can be consistently recovered from data as number of tasks increase. Specifically, each task has an associated ground-truth predictor $f_{t}^{\star}$, and we are interested in the covariance (second-moment) operator $\mathcal{S}$ that captures the common subspace spanned by these predictors. Since in practice we only observe finite samples per task and learn approximate predictors $\hat{f}_{t}$, two sources of error arise: (i) variability due to having finitely many tasks, and (ii) estimation noise within each task. Our goal is to establish conditions under which the empirical operators built from $\hat{f}_{t}$ concentrate around $\mathcal{S}$, and to show that the learned top-$k$ subspace converges to the true one, with convergence rates that separately reflect the number of tasks and the accuracy of per-task learning.
@@
Here $\mathcal{R}_{n_{t},D_{t}}(\mathcal{H})$ represents Rademacher complexity of the solutions within Hilbert space $\mathcal{H}$ over $n_{t}$ samples drawn i.i.d. from $D_{t}$ This form is satisfied, for example, by strongly convex regularized ERM in an RKHS (e.g., kernel ridge regression or NTK ridge), under bounded kernel norm and sub-Gaussian response noise (Bartlett & Mendelson, 2003).
@@ 3.1 Analysis methodology
Since there is no current method that enables us to compare subspaces of models with different architectures, we focus on large number of models trained on the same architecture. To this end, we perform analysis using Low rank adapters (Hu et al., 2021) (LoRA) as well as classical weights of transformer and CNN (Convolutional Neural Network) architectures. For all our experiments, unless stated otherwise, we perform Order 1-2 HOSVD only, to ensure that our methodology works even in the simplest case. Algorithm˜1 provides the algorithm we implement. Refer to Appendix˜B for discussion regarding secondary subspace and how to choose the number of top components.
@@ 3.2.1 Lower-rank joint subspaces in CNNs, LoRA and Finetuned models
In smaller and conventional architectures such as CNNs, evidence for universal structure has been more limited but suggestive. Early work observed that the first convolutional layer often learns Gabor-like filters across diverse vision tasks (Krizhevsky et al., 2012). More recently, works report recurring eigenvectors in certain CNN layers trained on natural images (Guth et al., 2024; Guth & Ménard, 2024).
@@
In order to conduct a more real-world experiment, we choose to run the subspace analysis for LoRA Hu et al. (2021) models simply because they are available in abundance in public domain. Given LoRA models distinctly capture task specific directions as they show weak alignment with the original weights Hu et al. (2021), they form a good main model parameter alternative to run our subspace analysis and verify whether this holds true. We spectrally decompose (Section 3.1) LoRA’s submatrices individually, each concatenated across all the available finetuned LoRAs and choose top $k$ spectral basis. This setup allows us to truly stress test the Universal Subspace.
@@
In our first experiment, we use LoRA models trained on 500 natural instruction tasks (Wang et al., 2022) using Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) as the base (Brüel-Gabrielsson et al., 2024). Each LoRA, individually, is at least of rank 16. Figure 3 shows the results of our analysis. 3(a) shows the subspace analysis of individual layers of the Mistral model. Each bar corresponds to the Eigenvalue or explained variance of every unique component (unique for all models, trained in disjoint/different datasets), showing that the parameters of all 500 models can be well approximated by a finite low-rank subspace for all layers. Note that for visualization purposes, the graphs only show the top 100 components, but there exists a significantly larger number of basis. 3(b) provides a summary of this subspace analysis, (summarized) for all of the layers of the 500 Mistral models, further validating the idea of a layer-wise universal subspace. Please refer to the appendix for all other and larger versions of the given spectral plots.
@@
We extend our analysis to text-to-image generation using Stable Diffusion-XL (Podell et al., 2023). A universal subspace is extracted from publicly available LoRAs on HuggingFace (von Platen et al., 2022). When projecting individual LoRAs into this subspace, the resulting generations preserve visual quality and style (Figure 5). CLIP-based evaluations (Table 1) show that the universal subspace even outperforms individual LoRAs in some cases, possibly due to denoising effects previously observed in (Sharma et al., 2023).
@@
In order to test the ability of condensing many models into a single universal subspace, we compare our method with SOTA model merging/combination methods in Table 2. We compare our universal subspace inspired combination approach against six state-of-the-art, gradient-free baselines: RegMean (Jin et al., 2023), Task Arithmetic (TA) (Ilharco et al., 2023), TIES (Yadav et al., 2023), DARE-TIES (Yu et al., 2024), KnOTS-TIES, and KnOTS-DARE-TIES (Stoica et al., 2025). RegMean aligns task-specific updates by solving a layer-wise linear regression problem, requiring transformation matrices for each model. TA merges models by linearly combining parameters, but relies on tuning scaling coefficients on a validation set for optimal performance. TIES extends TA with magnitude-based pruning and sign conflict resolution, introducing additional hyperparameters such as pruning thresholds, while DARE-TIES combines random Bernoulli pruning with TIES’ sign resolution, also requiring tuning of pruning probability. KnOTS-TIES and KnOTS-DARE-TIES further apply SVD-based subspace alignment before merging, but still inherit the need for coefficient or pruning hyperparameter selection. In contrast, our universal subspace method, analytically computes the merging coefficients based solely on the geometry of a shared, low-rank universal subspace identified across models, requiring no iterative tuning or validation data-although optional finetuning is possible if data is available. Furthermore, because our subspace is intrinsically low-rank, the merged model contains significantly fewer parameters than any individual models, offering both computational efficiency and theoretical alignment guarantees not present in the baselines. Empirically, our approach achieves higher average accuracy (see Table 2), while reducing parameter count, thus enabling scalable and robust model merging without heuristic pruning or validation overhead. We note that we did not optimize our merging process and better results nearing finetuned performance may be achieved.
@@ 3.3 Low rank shared universal subspaces in classical weights
We further extend this analysis to 50 finetuned LLaMA3-8B models, 177 GPT-2 models, and Flan-T5 models (trained on GLUE Wang et al. (2019) datasets) again sourced from HuggingFace without filtering. As shown in Figure 6, a small number of directions capture dominant structure across models spanning diverse and distinct datasets and tasks. More details are provided in the Section˜B.2. This is, to our knowledge, the first instance of compressing such a large and diverse collection of foundation models into a unified subspace, highlighting its potential for large-scale model reuse and environmental efficiency.
@@ 3.3.1 Finding universal subspaces and applying them to future tasks
We present two experiments - Image Classification using ViT-base and Natural Language Understanding using GLUE benchmark Wang et al. (2019) with RoBERTa-base model. Both involve creating a universal subspace using publicly available LoRA adapters. Details are provided in the Appendix˜C. For the GLUE benchmark, we follow the same setup as  (Kopiczko et al., 2023) considering the 6 tasks - CoLA, MRPC, SST-2, QNLI, RTE and STS-B while omitting the time-intensive MNLI and QQP tasks. We initialize our universal subspace using a leave-one-out-setup, where the subspace is calculated using components of all but one LoRA adapter for which the coefficients are learned. For image classification, we utilize publicly available ViT LoRAs to extract our universal subspaces taking care that the data any of these pretrained LoRAs have not seen the data we will be training our coefficients on.
@@ 4 Discussion
Why do these universal subspaces emerge? While the precise mechanisms driving this phenomenon remain an open area of investigation, several theoretical factors likely contribute to the emergence of these shared structures. First, neural networks are known to exhibit a spectral bias toward low-frequency functions, creating a polynomial decay in eigenvalues that concentrates learning dynamics into a small number of dominant directions (Belfer et al., 2024; Bietti et al., 2019). Second, modern architectures impose strong inductive biases that constrain the solution space: convolutional structures inherently favor local, Gabor-like patterns (Krizhevsky et al., 2012; Guth et al., 2024), while attention mechanisms prioritize recurring relational circuits (Olah et al., 2020; Chughtai et al., 2023). Third, the ubiquity of gradient-based optimization – governed by kernels that are largely invariant to task specifics in the infinite-width limit (Jacot et al., 2018) – inherently prefers smooth solutions, channeling diverse learning trajectories toward shared geometric manifolds (Garipov et al., 2018). If these hypotheses hold, the universal subspace likely captures fundamental computational patterns that transcend specific tasks, potentially explaining the efficacy of transfer learning and why diverse problems often benefit from similar architectural modifications.
@@ 5 Limitations and Future Work
Although we provide conclusive results towards the existence and utility of universal shared subspaces, the current analysis has scope for future research, such as limited interpretability of the shared subspace and the corresponding directions. While it is a critical area of research, it is extremely cumbersome to demonstrate interpretability of the principal directions for each layer of the network. To the best of our knowledge we are not aware of any other literature that performs such an in-depth analysis of the weight space of large models across diverse tasks, data modalities and model architectures. The current approach to approximating a universal subspace relies on pretrained task-specific models (predictors) for tasks, which may not be readily available for new tasks. An interesting direction for future research would be to explore model independent methods for learning a universal shared subspace, potentially derived directly from data. Furthermore, the conditions proposed in Ortiz-Jimenez et al. (2023) for enabling task arithmetic rely on localized eigenfunctions which are not conducive to learning a shared universal subspace. As a result, performing task arithmetic within the current framework of a shared universal subspace is non-trivial and warrants further investigation.